OpenStudy (anonymous):

Find the values of the trigonometric functions of theta from the given information sec(theta)= 2 , sin(theta) < 0

OpenStudy (anonymous):

\[\huge \text{sec}(\theta)=2, \text{sin}(\theta)<0\] \[sec(\theta)=1/cos(\theta) \]

OpenStudy (campbell_st):

so 1/cos = 2 or 2cos = 1 or cos = 1/2 so you have 4th quadrant... cos is positive and sin is negative so the angle is 360 - th \[\cos^{-1}(1/2) = 60\] so \[\theta = 300\]